two bodies of mass 40 kg and 80kg are at a distance of 0.15 m from each other the force of gravitation between the bodies is 1.0 mg wt calculate G if g =10m/ s^2

Grade:11

1 Answers

Arun
6 years ago
Dear student
 
As we know
 
F = G* m1* m2/r²
 
Now F = (1/1000)* 9.81 N = 0.00981 N
 
Put all the values
You will simply find G.
 
G = 7*10-11 N-m²/kg²
 
Regards
Arun (askIITians forum expert)